OpenStudy (anonymous):

How much water must be evaporated from 150 liters of 4% salt solution to obtain a solution that is 6% salt?

hero (hero):

 150*.04 - 0x = .06y 150 - x = y 6 = .06y 150 - y = x y = 6/.06 y = 100 150 - 100 = x 50 = x